OpenStudy (anonymous):

-1, 2 ,12, 40 what will be the next term? please reply quick

OpenStudy (amistre64):

whats the rush?

OpenStudy (amistre64):

0,3,13,41 if we shift it but i dont see a pattern in it at the moment

OpenStudy (amistre64):

3 10 28 7 18 9 not enough terms to determine it that way either

OpenStudy (nenadmatematika):

-1+1=0 (zero at 40)....2+2=4(4 at 40)...maybe that's some kind of the pattern...I guess not...

OpenStudy (anonymous):

did they give u any option on this question

OpenStudy (anonymous):

@asesde No

OpenStudy (anonymous):

I think it would be 49

OpenStudy (asnaseer):

There are lots of patterns you could fit to this sequence. One pattern is as follows:\[a_n=(2n-1)\times2^n\]for \(n=0,1,2...,\infty\)